# MtgrBoost: Boosting Large-scale Generative Recommendation Models in Meituan

### v1

@@ Abstract
Recommendation is crucial for both user experience and company revenue, and generative recommendation models (GRMs) are shown to produce quality recommendations recently. However, existing systems are limited by insufficient functionality support and inefficient implementations for training GRMs in industrial scenarios. As such, we introduce MTGRBoost as an efficient and scalable system for GRM training. Specifically, to handle the real-time insert/delete of sparse embedding entries, MTGRBoost employs dynamic hash tables to replace static tables. To improve efficiency, MTGRBoost conducts dynamic sequence balancing to address the computation load imbalances among GPUs and adopts embedding ID deduplication alongside automatic table merging to accelerate embedding lookup. MTGRBoost also incorporates implementation optimizations including checkpoint resuming, mixed precision training, gradient accumulation, and operator fusion. Extensive experiments show that MTGRBoost improves training throughput by 1.6×1.6$\times$1.6 ×2.4×2.4$\times$2.4 × while achieving good scalability when running over 100 GPUs. MTGRBoost has been deployed for many applications in Meituan and is now handling hundreds of millions of requests on a daily basis.
1 Introduction
@@
As a leading e-commerce platform for lifestyle services, Meituan spans multiple business verticals such as delivery, travel, and health. In 2024, we served over 770 million transacting users and processed more than 98 million orders at the daily peak. As shown in Figure 1, recommendation models [36, 47, 25, 13, 45] are at the core of our business by effectively filtering information and personalizing services for users. High-quality recommendations are crucial to sustain high customer retention rates, ensure merchant satisfaction, and enhance company revenue. Currently, our recommendation model incorporates terabytes of user behavioral data and item metadata for training on a daily basis and produces billions of predictions for various services.
@@
Following the advancements of generative language models [1, 26, 33, 34], generative recommendation models (GRMs) [44, 4, 37, 43, 11, 8, 42] have gained significant attention from both academia and industry. In particular, a GRM consists of two modules, i.e., sparse embedding tables that map discrete categorical features to continuous embedding vectors, and a Transformer-based dense model that predicts the next user action by modeling the action sequences of users. In contrast, earlier deep recommendation models (DRMs) [36, 47, 25, 41] utilize multilayer perceptrons to capture the interaction between user and each individual action. As shown in Figure 2, GRM achieves superior accuracy compared with DRM, primarily due to its ability to model the entire action sequence of each user with self-attention. However, GRM has greater computation complexity due to its attention mechanism, which scales quadratically with sequence length, while DRM scales only linearly. As such, to enjoy the good accuracy of GRMs and keep up with the high speed of training data generation, Meituan requires an efficient and scalable system for training GRMs.
Figure 1: Applications of recommendation model in Meituan.
Extensive researches have been conducted to optimize the execution of Transformer-based dense models [3, 30, 28, 24, 9]. For instance, FlashAttention [3] accelerates attention computation by materializing less intermediate results and hence reducing the accesses GPU global memory. Distributed training frameworks such as Megatron [30] and DeepSpeed [28] enhance scalability by using hybrid parallel training strategies that combine data, model, and pipeline parallelism. These developments motivate a migration from TensorFlow [31] to PyTorch [15] for industrial recommendation systems. In particular, TorchRec [10] is the state-of-the-art PyTorch-based system for recommendation model training, and we develop MTGRBoost based on it by integrating efficient attention computation kernels.
@@
However, we identify two fundamental inefficiencies in TorchRec when handling the sparse embedding tables of GRMs. ❶ TorchRec’s static embedding tables can not process dynamic embedding entries and table capacity expansion. Specifically, additional embeddings cannot be allocated to new users and items (e.g., merchants updating menus) at real-time from fixed-size static tables. Although some default embeddings may be used in this case, model accuracy will be degraded. Static tables inherently lack dynamic expansion capabilities. Thus, static tables typically require pre-allocation of capacity exceeding actual requirements to prevent ID overflow, inevitably leading to inefficient memory utilization. ❷ For embedding lookup, TorchRec can fetch the same embedding entry multiple times due to duplicate feature IDs in sequences, which causes a long communication time and wastes network bandwidth. Moreover, TorchRec requires labor-intensive manual configuration to merge the embedding tables, which is a common and effective operation to accelerate training. The two limitations severely constrain GRM’s practical deployment and compromise training efficiency.
@@
Figure 2: AUC (for accuracy, higher the better) and model complexity for DRM and GRM. An accuracy improvement of even 0.1% is crucial for practical recommendation.
To tackle the challenges, MTGRBoost features a hash-based dynamic embedding table and two associated optimizations, i.e., automatic table merging and lookup deduplication. The hash-based embedding table eliminates fixed-capacity constraints by decoupling the storage of keys (i.e., embedding index) and values (i.e., embedding vector) and uses hash functions to map arbitrary feature IDs to embedding index. We tailor the capacity expansion process by maintaining two value allocations and only doubling the key storage. The insight is that the network communication for migrating the keys is significantly lower than moving the high-dimensional embedding vectors as for the static tables. Moreover, we adopt two-stage ID unique operation to reduce redundant IDs before and after ID communication to avoid repetitive embedding transfers among the devices. We also design a feature configuration interface to implement automatic table merging, which can reduce the number of embedding lookup operators and thus speed up the process.
Besides the sparse embeddings, MTGRBoost incorporates a suite of additional optimizations to implement an efficient GRM training system. We observe that variable-length sequences cause significant load imbalances among the GPUs. Since retaining complete sequences is crucial for accuracy, we cannot truncate or pad sequences for load balancing as done for LLMs. Instead, we propose a light-weight dynamic sequence balancing technique, which dynamically adjusts the batch sizes for GPUs such that the GPUs have similar workload. Moreover, MTGRBoost supports checkpoint resumption, mixed-precision training, gradient accumulation, and operator fusion. In particular, MTGRBoost allows saving and loading checkpoints using different numbers of GPUs, uses different precisions for the dense model, hot embeddings, and cold embeddings to reduce computation, conducts efficient sparse gradient accumulation through sparse aggregation, and accelerates forward computation by fusing multiple kernels in the attention modules.
We evaluate MTGRBoost and its designs using 300 million real user sequences generated during a week at Meituan. Scalability experiments show that MTGRBoost achieves close to linear speedup in training throughput when using more GPUs. Compared with TorchRec, MTGRBoost reduces the training time by 2.44×2.$4$. Ablation experiments show that our designs are effective. In particular, dynamic sequence balancing achieves near-optimal load balance and improves system throughput by 1.75$\times$ compared to the baseline. Moreover, automatic table merging and lookup de-duplication significantly reduce network communication, yielding an improvement of 53% in throughput.
@@
    We present MTGRBoost as an efficient and scalable system for training generative recommendation models (GRMs) at Meituan in real production environment.
@@
    We observe that existing systems are inefficient when handling the sparse embeddings of GRMs. As such, we propose techniques including dynamic table to adapt changing features, automatic table merging to fuse operators, and lookup de-duplication to avoid repetitive communication.
@@
    We build MTGRBoost as a complete and efficient system for GRM training with optimizations including load balancing, flexible checkpoint resumption, mixed-precision computation, gradient accumulation, and kernel fusion.
@@
    We conduct extensive experiments on real data from production to evaluate MTGRBoost. The results show that MTGRBoost is efficient and scalable, and our designs are effective in improving system efficiency.
Figure 3: Model architecture of Meituan’s GRM.
2 GRM in Meituan
Model Architecture. Figure 3 illustrates the model architecture of the generative recommendation model (GRM) in Meituan, which consists of four core components: input sequence, embedding layer, hierarchical sequential transduction units (HSTU) layer, and multi-gate mixture-of-experts (MMoE) layer. In particular, the input sequence $\mathbf{T}=[\mathrm{T}_{con},\mathrm{T}_{hst},\mathrm{T}_{exp}]$ includes contextual sequences (i.e., user features), historical sequences (e.g., click and purchase actions), and exposed sequences (i.e., real-time actions). Since the deep model cannot directly process discrete category data, the embedding layer $\phi_{emb}$ is used to convert the feature IDs into continuous embedding vectors $\mathbf{E}=\phi_{emb}(\mathbf{T})\in\mathbb{R}^{F\times d}$, where $F$ and $d$ denote the number of features and embedding dimensions, respectively.
Subsequently, we employ multiple HSTU layers – an attention-based Transformer architecture – to learn user sequence embeddings through user and item interactions. As shown in Figure 3, each HSTU is equipped with four attention-based sub-layers for user feature extraction $\mathbf{U}$, item feature extraction $\mathbf{O}$, and feature interaction $\mathbf{H}$. We formulate the calculation process as follows:
$\mathbf{U,Q,K,V}=\mathrm{Split}(\phi_{1}(\mathrm{MLP}(\mathbf{E}))),$ (1)
$\mathbf{O}=\phi_{2}(\mathbf{Q}\mathbf{k}^{T})\mathbf{V},$ (2)
$\mathbf{H}=\mathrm{MLP}(\mathrm{Norm}(\mathbf{O}\odot\mathbf{U})),$ (3)
where $\phi_{1}$ and $\phi_{2}$ are the SiLU activation function. MMoE directs embeddings $\mathbf{H}$ from the HSTU layer to multiple expert models. Each expert model is equipped with a gating network that learns the model’s weights. We finally aggregate the output embeddings of the top-$k$ expert models. That is:
$y=\sum_{i=1}^{k}g_{i}(\mathbf{H})\cdot\mathrm{Expert}_{i}(\mathbf{H}).$ (4)
We use cross entropy loss to optimize click-through rate and conversion rate.
Different from GRM and DRM. Our model builds upon GR [44] by introducing a novel prediction head enhanced through MMoE. This design replaces GR’s MLP with MMoE, which employs task-specific gate networks to dynamically balance multiple objectives, improving accuracy in multi-task scenarios. Compared to MLP’s parameter-intensive tuning, MMoE reduces computational overhead by selectively activating only the top-k expert outputs. Additionally, our MTGRBoost diverges from DRM in batch construction. As illustrated in Figure 4, DLRM’s pairwise batches contain redundant user computations (e.g., "User 1" features are processed twice). In contrast, MTGRBoost employs a sequence-wise approach that consolidates multiple user interactions into a single sample, which ensures each user feature is computed once. Notably, user features contain orders of magnitude more tokens than item features (e.g., 10,000 vs. 100), enabling MTGRBoost to significantly accelerate training and improve scalability.
Figure 4: Batch structure of DRM and GRM. “U” denotes users, “P” denotes products. DRM and GRM use a pairwise and sequence-wise batch construction method, respectively.
3 MTGRBoost System Overview
MTGRBoost facilitates the distributed training of GRMs across multiple devices. For parallelism, we follow the TorchRec’s design and employ a hybrid strategy: model parallelism for sparse models and data parallelism for dense models. This combined approach is informed by two considerations. Firstly, data parallelism is impractical for the large embedding table, as no single device can accommodate the entire table. Secondly, the dense models are relatively small, which renders data parallelism more suitable. MTGRBoost is trained end-to-end with synchronous execution, ensuring model quality and accuracy.
Figure 5: The workflow of MTGRBoost for GRM training.
Workflow. We provide the detailed explanation as follow.
@@ 3 MTGRBoost System Overview
    Data I/O. As shown in Figure 5 (1), the training data is stored in partitioned Hive tables on HDFS, which utilizes a columnar storage format to optimize access speed and storage efficiency. Each column in the Hive table represents a feature. To maximize I/O throughput, the Hive tables are partitioned into smaller shards distributed across devices, which read data in parallel from their assigned shards. Additionally, we prefetch multiple next batches and overlap their loading with the computation of the current batch, thereby masking I/O latency.
@@
    Embedding Lookup. As shown in Figure 5 (2), the embedding lookup process employs All-to-all communication for cross-device embedding exchange. In particular, this process involves two data exchanges: First, devices transmit feature IDs requiring queries and receive corresponding IDs from peer devices through All-to-all communication. Subsequently, another All-to-all operation facilitates the exchange of retrieved embeddings between devices.
@@
    Forward Computation. The dense models within MTGRBoost, including HSTU and MMoE, have significantly fewer parameters compared to sparse model (i.e., embedding tables). As depicted in Figure 5 (3), data parallelism prepares a replication of the dense model for each device during forward propagation. Model parameters across all devices are consistently initialized by setting the same random seed. MTGRBoost partitions the total sequence batch into smaller micro-batches distributed to various devices, then performs forward computation in parallel.
  • $\bullet$
@@
    Backward Update. Figure 5 (4) illustrates the heterogeneous parameter update adopted by MTGRBoost in the hybrid parallel architecture. Sparse embedding tables aggregate shard gradients through model parallelism, where each device updates its locally stored embedding shards via All-to-all communication. Dense parameters (e.g., HSTU) employ All-Reduce communication for gradient synchronization before parameter update.
Pipeline. To further enhance training efficiency, we maximize parallelism through pipeline technology using three streams: copy, dispatch, and compute. Specifically, the copy stream loads sequences from CPU to GPU, the dispatch stream performs table lookups based on IDs, and the compute stream handles forward computation and backward updates. For instance, while the compute stream executes forward and backward passes for batch $\mathbf{T}$, the copy stream concurrently loads batch $\mathbf{T}+1$ to mask I/O latency. Upon completing backward updates for batch $\mathbf{T}$, the dispatch stream immediately initiates table lookups and communication for batch $\mathbf{T}+1$.
Figure 6: Dynamic embedding table in MTGRBoost. (a) Storage layout. We design a decoupled key-value storage architecture, where keys are stored in the Key Structure and values reside in a dedicated Embedding Structure. (b) Table lookup. For a given ID, we employ hash and probing functions to identify its index, and then retrieve the embedding’s memory address using an offset-based mechanism. (c) Capacity expansion. We prioritize key structure expansion and initiate embedding structure expansion only when the current memory chunk reaches capacity.
4 Efficient Processing for Sparse Embeddings
4.1 Dynamic Embedding Table
@@
In Meituan’s production environment, the recommendation system must dynamically adjust recommendations based on real-time user interactions. However, TorchRec cannot handle dynamic IDs due to its reliance on fixed-capacity static embedding tables. Static tables typically use default embeddings to provide embeddings for feature IDs that exceed the capacity, but it leads to model accuracy degradation. Furthermore, static table necessitates over-provisioning of initial capacity to anticipate potential expansion requirements, which results in inefficient memory utilization. These limitations motivated MTGRBoost’s implementation of dynamic embedding tables.
Storage Layout. MTGRBoost employs a decoupled architecture for hash table, which separates key and value storage into distinct structures. As shown in Figure 6 (a), the key structure maintains a lightweight mapping table containing keys and corresponding pointers to embedding vectors, while the embedding structure stores both the embedding vectors and auxiliary metadata (e.g., counters and timestamps) required for eviction policies like Least Recently Used and Least Frequently Used. This dual-structure design achieves two critical objectives: (1) enabling dynamic capacity expansion through replication of only the compact key structure rather than huge embedding data, and (2) optimizing key scanning efficiency by storing keys in compact memory layouts while accommodating potentially sparse key distributions.
Furthermore, we implement chunk-based allocation for the embedding structure. This design provides two benefits: first, it reduces memory fragmentation through bulk allocation of larger chunks rather than frequent small allocations, while enabling efficient single-operation deallocation; second, it enhances cache locality through sequential memory access patterns, thereby improving hash table access efficiency.
@@
Handling Hash Function and Collisions. MTGRBoost selects MurmurHash3 [2] as the hash function to determine where the embedding data is stored. MurmurHash3 processes input ID in 4-byte blocks through mixing operations (e.g., constant multiplication, bit rotation, XOR merging) to maximize entropy and ensure avalanche effects from single-bit changes. The multi-stage mixing generates uniformly distributed hash values through iterative nonlinear transformations.
@@
Hash collisions occur when different keys are mapped to the same index position by the hash function. Techniques such as chaining and open addressing are typically employed. MTGRBoost chooses open addressing to to resolve collisions due to its memory efficiency advantages over chaining. However, existing probing techniques exhibit vulnerability to clustering behavior under high load factors. To address this issues, we propose a novel grouped parallel probing technique. This approach begins by ensuring that the base step size $S$ is an odd number through a bitwise OR operation with 1, thereby guaranteeing that all positions in the table can be accessed during probing. Next, we associate the step size with the key to prevent clustering that results from overlapping probing sequences of different keys. Finally, we multiply the step size by the thread group number, enabling distinct thread groups to probe different regions of the table. Formally, this technique is expressed as follows:
$S=(k\%(M/\mathsf{threads}-1)+1\mid 1)*\mathsf{threads},$ (5)
@@
where $k$ refers to the key’s value, $M$ represents the hash table size, $\mathsf{threads}$ indicates the number of thread groups, and $\mid$ denotes the bitwise OR operation. Grouped parallel probing leverages the highly parallel capabilities of GPUs, ensuring a uniform distribution of the probing sequence. Moreover, we use coprime conditions to prove that our grouped parallel probing technique can traverse all slots in the hash table. The proof is provided in Appendix A.
Theorem 1
@@
$\forall i,j\in\left[0,M-1\right],i\neq j\Rightarrow h_{i}\neq h_{j}.$ (6)
@@
Table Lookup. With the hash function and probing method in place, we can now implement dynamic table lookup effectively. As shown in Figure 6 (b), a hash value is computed for a given user ID, and probing identifies the index corresponding to this ID within the hash table (steps 1-3). Using the starting address of the key structure in memory, along with the index and pointer offset, we determine the embedding pointer $p$ for the user ID (step 4), formally defined as follows:
$p\!=\!st\_add\!+\!index*row\_offset\!+\!pointer\_offset,$ (7)
where $st\_add$ denotes the start address of key structure,and $row\_offset=key\_offset+pointer\_offset$. Finally, we use this pointer, along with the embedding offset in the embedding structure, to extract the embedding vector (step 5). Unlike static tables, which cannot handle input data ranges exceeding their size, our dynamic table can accept changeable input value. Consequently, the dynamic table effectively addresses the limitation of static tables on managing dynamic IDs.
@@
Capacity Expansion. MTGRBoost initiates hash table expansion when slot occupancy reaches critical levels (typically at load factors >0.75). As depicted in Figure 6 (c), capacity follows a power-of-two progression, doubling iteratively until meeting system requirements. The migration process transfers the keys and pointers from the original key structure to the expanded version and then deallocating the legacy memory. Crucially, our design prioritize expanding the key structure while maintaining the chunk (i.e., embedding structure) capacity, thus optimizing expansion efficiency by avoiding bulk embedding data transfers. MTGRBoost maintains two chunks – a current chunk and a next chunk – to address embedding structure expansion requirements. Specifically, when the current chunk’s remaining capacity becomes insufficient to accommodate new embeddings, we store them in the next chunk. Concurrently, we retire the filled chunk and allocate a new chunk as the next chunk to preserve the dual-chunk configuration. This pre-allocation mechanism ensures immediate storage availability for new embeddings during expansion.
Figure 7: An example of offsets in static table and hash table.
4.2 Embedding Table Merging

### v2

# MTGenRec: An Efficient Distributed Training System for Generative Recommendation Models in Meituan

@@ Abstract.
Recommendation is crucial for both user experience and company revenue in Meituan as a leading lifestyle company, and generative recommendation models (GRMs) are shown to produce quality recommendations recently. However, existing systems are limited by insufficient functionality support and inefficient implementations for training GRMs in industrial scenarios. As such, we introduce MTGenRec as an efficient and scalable system for GRM training. Specifically, to handle real-time insertions/deletions of sparse embeddings, MTGenRec employs dynamic hash tables to replace static ones. To improve training efficiency, MTGenRec conducts dynamic sequence balancing to address the computation load imbalances among GPUs and adopts feature ID deduplication alongside automatic table merging to accelerate embedding lookup. Extensive experiments show that MTGenRec improves training throughput by $1.6\times$$2.4\times$ while achieving good scalability when running over 100 GPUs. MTGenRec has been deployed for many applications in Meituan and is now handling hundreds of millions of requests on a daily basis. On the delivery platform, we observe a 1.22% growth in user order volume and a 1.31% enhancement in online PV_CTR.
Recommendation Systems, Distributed Model Training
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NY
1. Introduction
@@
As a leading e-commerce platform for lifestyle services, Meituan spans multiple business verticals such as delivery, travel, and health. In 2024, Meituan served over 770 million transacting users and processed more than 98 million orders at the daily peak. As shown in Figure 1, recommendation systems (wang2017deep; zhou2018deep; naumov2019deep; jiang2019xdl; zhang2022picasso) are at the core of our business by effectively filtering history logs and personalizing services for users. High-quality recommendations are crucial to sustain high customer retention rates and enhance company revenue. Currently, our recommendation model utilizes terabytes of user action data and item metadata for training on a daily basis and produces billions of predictions for various services.
@@
Inspired by generative language models (achiam2023gpt; radford2018improving; touvron2023llama; vaswani2017attention), generative recommendation models (GRMs) (zhai2024actions; deng2025onerec; wang2023generative; yan2025unlocking; ji2024genrec; huang2025sessionrec; xu2025efficient) have gained significant attention from both academia and industry. Specifically, a GRM consists of two modules, i.e., sparse embedding tables that map discrete categorical features to continuous embedding vectors, and a Transformer-based dense model that predicts the next user action by modeling the action sequences of users. In contrast, earlier deep recommendation models (DRMs) (wang2017deep; zhou2018deep; naumov2019deep; xu2018deep) utilize multilayer perceptrons to capture the interactions between users and items. As shown in Figure 2, GRM achieves superior accuracy compared with DRM, primarily due to its ability to model the entire action sequence of each user with self-attention. However, GRM has greater computation complexity due to its attention mechanism, which scales quadratically with sequence length, while DRM scales only linearly. Thus, to enjoy the good accuracy of GRMs and keep up with the high speed of training data generation, Meituan requires an efficient and scalable system for training GRMs.
Figure 1. Recommendation workflow in Meituan.
@@
TorchRec (ivchenko2022torchrec) is a state-of-the-art PyTorch-based system for training GRMs. It enables efficient computation of attention scores (dao2022flashattention) and enhances scalability through hybrid parallel strategies, including data and model parallelisms (shoeybi2019megatron; rasley2020deepspeed; huang2019gpipe; narayanan2019pipedream). To leverage TorchRec’s benefits, we further develop our MTGenRec based on it. However, we identify two fundamental limitations in TorchRec that hinder its application in industrial recommendation systems. In particular, ❶ when TorchRec handles the sparse embedding tables of GRMs, its built-in static embedding tables cannot process dynamic embedding entries and table capacity expansion. Specifically, additional embeddings cannot be allocated to new users and items (e.g., merchants updating menus) in real-time from fixed-size static tables. Although default embeddings can be used in this case, model accuracy will be degraded. Moreover, static tables typically require pre-allocating more capacity than necessary to prevent overflow, leading to memory underutilization. ❷ For the embedding lookup process, TorchRec transmits same embedding entries multiple times due to duplicate embedding IDs in the user sequences, leading to a long communication time. Moreover, TorchRec requires labor-intensive manual configuration to merge embedding tables, which is a common and effective operation for accelerating lookups. The two limitations severely constrain the practical deployment of GRMs and compromise training efficiency.
@@
Figure 2. AUC (for accuracy, higher the better) and model complexity for DRM and GRM. Notably, an accuracy improvement of even 0.1% is crucial for practical recommendation.
To address these challenges, we propose a hash-based embedding table for changeable IDs and two-stage ID deduplication for lookup acceleration. The hash-based embedding table eliminates fixed-capacity constraints by mapping arbitrary feature IDs to embedding indices via a hash function. We decouple the storage of keys (i.e., embedding index) and values (i.e., embedding vector). This design prioritizes the expansion of the smaller key list over the larger value list, thereby enhancing the efficiency of table expansion. Additionally, we adopt a two-stage ID deduplication operation to eliminate repetitive IDs before and after ID communication, thus avoiding the repetitive transmission of embedding vectors among devices. We also design a table configuration interface for automatic table merging, which combines multiple embedding tables into one according to defined patterns (e.g., the same embedding dimension). This technique reduces the number of embedding lookup operators, and thus accelerating the process. Besides the sparse embeddings, we observe that variable-length sequences cause significant load imbalances among GPUs. Since retaining complete sequences is crucial for GRM’s accuracy, we cannot truncate or pad sequences as is done in LLMs. Instead, we propose a lightweight dynamic sequence batching technique that dynamically adjusts batch sizes to ensure a more balanced workload across GPUs.
[…]
@@
    We present MTGenRec as an efficient and scalable system for training generative recommendation models (GRMs) at Meituan.
@@
    To address the challenges caused by GRM’s sparse and dense models, we introduce dynamic embedding tables and ID deduplication for changeable feature IDs and lookup acceleration, and dynamic sequence batching for computation load balancing.
@@
    We conduct extensive experiments to evaluate MTGenRec. The results show that MTGenRec is efficient and scalable, and our designs are effective in improving system efficiency.
Figure 3. Model architecture of Meituan’s GRM.
2. MTGenRec System Overview
Model Architecture. Figure 3 illustrates the model architecture of the GRM in Meituan, which consists of four core components: input sequence, embedding layer, hierarchical sequential transduction units (zhai2024actions) (HSTU) layer, and multi-gate mixture-of-experts (ma2018modeling) (MMoE) layer. In particular, the input sequence $\mathbf{T}=[\mathrm{T}_{con},\mathrm{T}_{hst},\mathrm{T}_{exp}]$ includes contextual sequences $\mathrm{T}_{con}$ (i.e., user features), historical sequences $\mathrm{T}_{hst}$ (e.g., click and purchase actions), and exposition sequences $\mathrm{T}_{exp}$ (i.e., candidate items). Since the model cannot directly process discrete category data, the embedding layer $\phi_{emb}$ is used to convert the feature IDs into continuous embedding vectors $\mathbf{E}=\phi_{emb}(\mathbf{T})\in\mathbb{R}^{F\times d}$, where $F$ and $d$ denote the number of features and embedding dimension, respectively.
Subsequently, we employ multiple HSTU layers – a self-attention based Transformer (vaswani2017attention) architecture – to learn user sequence embeddings from item interactions. As shown in Figure 3, each HSTU is equipped with four attention-based sub-layers for user feature extraction $\mathbf{U}$, item feature extraction $\mathbf{O}$, and feature interaction $\mathbf{H}$. The computation process is expressed as follows:
(1) $\mathbf{U,Q,K,V}=\mathrm{Split}(\phi_{1}(\mathrm{MLP}(\mathbf{E}))),$
(2) $\mathbf{O}=\phi_{2}(\mathbf{Q}\mathbf{k}^{T})\mathbf{V},$
(3) $\mathbf{H}=\mathrm{MLP}(\mathrm{Norm}(\mathbf{O}\odot\mathbf{U})),$
where $\phi_{1}$ and $\phi_{2}$ are the SiLU (elfwing2018sigmoid) activation function. MMoE directs embeddings $\mathbf{H}$ from the HSTU layer to multiple expert models. Each expert model is equipped with a gating network that learns the model’s weights. We finally aggregate the output embeddings of the top-$k$ expert models. That is:
(4) $y=\sum_{i=1}^{k}g_{i}(\mathbf{H})\cdot\mathrm{Expert}_{i}(\mathbf{H}).$
We select the cross entropy loss as the training objective.
Distributed Training. MTGenRec facilitates the distributed training of GRMs across multiple devices. We follow TorchRec’s design and employ a hybrid parallel strategy: model parallelism for sparse models and data parallelism for dense models. This design is driven by two considerations. First, data parallelism is impractical for the large embedding table, as no single device can store the entire embedding table. Second, dense models are relatively small, making data parallelism more suitable. MTGenRec is trained end-to-end with synchronous execution to ensure model quality and accuracy. To further enhance training efficiency, we introduce a pipeline that preprocesses the next data batch concurrently with the model computation. We detail the pipeline design in Appendix B.1. Next, We provide an introduction to the workflow of MTGenRec as follow.
Figure 4. The workflow of MTGenRec for GRM training.
@@ 2. MTGenRec System Overview
    Data I/O. As shown in Figure 4, the training data is stored in partitioned Hive tables on HDFS, which utilizes a columnar storage format to optimize access speed and storage efficiency. Unlike DRM that pairs a user with a single item in a data sample, GRM uses a sequence-wise approach that pairs a user with multiple items. This design avoids the repeated computation of a user feature. We provide an explanation in Appendix A.
@@
    Embedding Lookup. The embedding lookup uses All-to-all communication for cross-device embedding exchange. In particular, this process involves two data exchanges: First, devices transmit the required feature IDs and receive corresponding IDs from peer devices. Subsequently, a collection operation facilitates the exchange of retrieved embeddings between devices.
@@
    Forward Computation. After embedding lookup, we encode the input sequence through the dense model (i.e., HSTU and MMoE layers) and then calculate the loss function. Model parameters across all devices are consistently initialized by setting the same hyperparameter and random seed.
@@
    Backward Update. MTGenRec employs a heterogeneous parameter update strategy in the backward propagation. Specifically, sparse embedding tables aggregate gradients through model parallelism, where each device updates its locally stored embedding shards. Dense parameters employ All-Reduce communication for gradient synchronization before parameter update.
3. Core Designs for MTGenRec
3.1. Dynamic Embedding Table
@@
In a real production environment, the recommendation system must handle real-time insertions and deletions of sparse embedding entries (e.g., merchants adding or removing products). However, TorchRec cannot handle dynamic entries due to its static embedding tables (table capacity is fixed), and it typically uses default embeddings for feature IDs that exceed the table capacity. However, this method may lead to model accuracy degradation. Additionally, static tables require over-provisioning of initial capacity to anticipate potential expansion requirements, which results in inefficient memory utilization. These limitations motivate MTGenRec’s implementation of hash-based embedding tables.
Figure 5. Dynamic embedding table in MTGenRec.
Storage Layout. MTGenRec employs a decoupled hash table architecture to separate keys and values into distinct structures. As shown in Figure 5 (a), the lightweight key structure maintains a mapping table of keys and pointers to embeddings, while the embedding structure stores embeddings and auxiliary data (e.g., timestamps). This design allows for efficient table expansion by prioritizing the expansion of the smaller key structure over the larger embedding structure. Additionally, we implement chunk-based allocation for the embedding structure. Specifically, multiple embeddings are stored within a single chunk, and insertions/deletions are managed on a per-chunk basis. This approach enhances efficiency, as it requires only a single chunk allocation or deletion operation, rather than multiple individual operations.
@@
Hash Function and Collision Handling. MTGenRec selects MurmurHash3 (murmurhash) as the hash function to determine key indices due to its efficient hash value calculation. MurmurHash3 processes input IDs in 4-byte blocks through mixing operations, which maximize entropy and ensure avalanche effects from single-bit changes. The multi-stage mixing generates uniformly distributed hash values through iterative nonlinear transformations. Hash collisions occur when different keys are mapped to the same index by a hash function. Common techniques for handling collisions include chaining and open addressing. We employ open addressing due to its memory efficiency advantages over chaining. However, existing probing techniques are vulnerable to clustering under high load factors (bender2022linear).
@@
To address this issue, we propose a novel grouped parallel probing technique. This approach first ensures that the base step size $S$ is an odd number by performing a bitwise OR operation with 1, guaranteeing that all positions in the table can be accessed during probing. Next, we associate the step size with the key to prevent clustering caused by overlapping probing sequences of different keys. Finally, we multiply the step size by the thread group number, enabling distinct thread groups to probe different regions of the table. Formally, the computation is expressed as follows:
(5) $S=(k\%(M/\mathsf{threads}-1)+1\mid 1)*\mathsf{threads},$
@@
where $k$ refers to the key’s value, $M$ represents the hash table size, $\mathsf{threads}$ indicates the number of thread groups, and $\mid$ denotes the bitwise OR operation. We use coprime conditions to prove that our grouped parallel probing technique can traverse all slots in the hash table. The proof is provided in Appendix B.2.
Theorem 3.1.
@@
(6) $\forall i,j\in\left[0,M-1\right],i\neq j\Rightarrow h_{i}\neq h_{j}.$
@@
Table Lookup. With the hash function and probing method in place, we can now implement dynamic table lookup. As shown in Figure 5 (b), a hash value is computed for a given user ID, and probing identifies the index corresponding to this ID within the hash table (steps 1-3). Using the starting address of the key structure in memory, along with the index and pointer offset, we determine the embedding pointer $p$ for the user ID (step 4):
(7) p=st_add+indexrow_offset+pointer_offset,p\!=\!st\_add\!+\!index*row\_offset\!+\!pointer\_offset,
where st_addst\_add denotes the start address of the key structure. Finally, we use this pointer, along with the embedding offset in the embedding structure, to extract the embedding vector (step 5).
@@
Capacity Expansion. When the slots in the hash table approach full capacity, expansion becomes necessary. MTGenRec triggers this expansion when the slot occupancy reaches critical levels (i.e., load factor ¿ 0.75). As shown in Figure 5 (c), the capacity increases in a power-of-two progression, doubling iteratively to meet the required capacity. The migration process transfers the original key structure to the expanded version, and the legacy memory is deallocated. Notably, our design prioritizes expanding the key structure while maintaining the chunk (i.e., embedding structure) capacity, thus optimizing expansion efficiency by avoiding bulk transfers of embedding data. To manage the expansion of the embedding structure, MTGenRec maintains two chunks: a current chunk and a next chunk. When the current chunk’s remaining capacity is insufficient for new embeddings, they are stored in the next chunk. Concurrently, we retire the filled chunk and allocate a new chunk as the next chunk to preserve the dual-chunk configuration. This pre-allocation mechanism ensures immediate storage availability for new embeddings during the expansion process.
3.2. Embedding Table Merging